OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form.

OpenStudy (anonymous):

helpp pleaseeeeee AAAAASSSSAAAAPPPPP

OpenStudy (anonymous):

Find the slope first... (6-(-3))/(3-0) which is 3. Then use point slope form: y-y1=m(x-x1) and plug in 3 for m (slope) and use either point for x1,y1. Like this: y-6=3(x-3). Then simplify to get slope intercept form: y-6=3x-9 y=3x-3.

OpenStudy (anonymous):

all right check the way.. (3,6) and (0,3) from slope = Y2- Y1/X2 -x1 where by -3 -6/0 - 3 -9/-3 = 3 slope is 3 from slope =Y - Y1/X - X1 3 =y -6/x-3 y-6 =3(x-3) y-6=3x-9 y-6 = 3x -(9+6) y=3x -3 The equation is y = 3x - 3 Got it am Joel